Fill in the blanks.
 (b) 13 + _____ = 0
(c) 12 + (–12) = ____
(d) (–4) + ____ = –12
 
iaemfuewgnuersiojmhnbiorjhmneru

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(b) 13 + (-13) = 0

(c) 12 + (-12) = 0

(d) (-4) + (-8) = -12


​Thanks